# Image Corruption-Inspired Membership Inference Attacks against Large Vision-Language Models

### v1

@@ 2.1 Large Vision-Language Models
Large Vision-Language Models (Bai et al., 2025; Chen et al., 2024a, b; Liu et al., 2024b, a; Chen et al., 2024c; Tong et al., 2024; Zhu et al., 2025; Young et al., 2024; Zhu et al., 2023; Chen et al., 2023; Li et al., 2024c; Du et al., 2025), also known as multimodal large language models (Fu et al., 2024) are developing rapidly due to the success of language models (Wang et al., 2024; Zhao et al., 2023; Chiang et al., 2023; Touvron et al., 2023a, b; Grattafiori et al., 2024; Yang et al., 2025, 2024; Jiang et al., 2023; Team, 2023; Penedo et al., 2023) such as Qwen (Bai et al., 2023a; Yang et al., 2024) and LLaMA series (Touvron et al., 2023a, b). These models, like LLaVA 1.5 (Liu et al., 2024c), can generate textual outputs given textual questions and images.
@@ 2.2 Membership Inference Attack
With the development of multimodal learning, there are also some works exploring MIA on multimodal models (Hu et al., 2022b; Ko et al., 2023; Li et al., 2024b). We focus on vision-text here. Liu et al. (2021) studies MIAs on image encoder models such as CLIP vision encoder (Radford et al., 2021). It calculates the similarity scores between the augmented images’ embeddings and the scores are then used to train a classifier to infer the member status of an image. Ko et al. (2023) work on detecting whether an image-text pair is in the training data of CLIP models (Radford et al., 2021). Li et al. (2024b) investigates the single-modality MIA in LVLMs, which is a more practical scenario. They calculate the Rényi Entropy (Rényi, 1961) based on different slices of logits. Hu et al. (2025) found that member data and non-member data have different sensitivity to temperature. They then perform four attack methods under four different settings to detect whether a group of images is used in the training stage.
@@ 3.3 Intuition for Performing MIA
As we are only given the image, how to fully utilize the image to determine if the image remains a challenging question. Inspired by Liu et al. (2021) that a CLIP vision encoder tends to overfit to its training data and member images show higher similarity for two augmented versions, we assume that the image embedding of a member image from the vision encoder and modality alignment module should be more robust to image corruption than that of non-member images. The intuition is, if the model has seen the image and memorized the image, then even if the image is corrupted, e.g., some details missing, the model can still recall the details and result in embedding similar to the original one. To verify our assumption, for each image, we first apply Gaussian Blur (Bradski, 2000) as the corruption method to obtain the corrupted image, where Gaussian Blur is a smoothing technique that reduces image detail and noise by averaging pixels with a weighted Gaussian kernel. We compute the similarity score between the embedding of the original image and that of its corrupted version. The chosen model is LLaVA-1.5-7B (Liu et al., 2024a), and the dataset is the Img_Flickr dataset constructed by Li et al. (2024b). The similarity scores of member images and non-member images are shown in Figure 3. We can observe that member images normally show a higher similarity score than non-member images, which aligns well with our assumption. The scores show a clear difference between the two groups, which suggests that the similarity score between the original image embedding and its corrupted version’s embedding can be used to perform an image membership inference attack. Thus, we can utilize the robustness of member image embedding to determine its membership.
@@ 4.1 White-Box MIA via Image Similarity
where $\mathbf{e}_{i}^{v}$ is the embedding of the $i$-patch of image $\mathbf{x}_{t}$ obtained by Eq. 1 and $\mathbf{e}_{i}^{v^{\prime}}$ denotes the $i$-patch embedding of the corrupted image $\mathbf{x}_{t}^{\prime}$. $N$ is the number of patches. Sim is a similarity function. We use cosine similarity here. The larger the $s_{img}$ is, the more likely the image is viewed as a member image. The target image $\mathbf{x}_{t}$ is predicted as a member image if $s_{img}$ is bigger than a threshold $\lambda$.
Algorithm 1 summarizes the image similarity-based attack.
@@ 4.2 Black-box MIA via Text Similarity
where $Emb()$ denotes a text embedding model, such as OpenAI’s text-embedding-3-small model [^1] https://platform.openai.com/docs/models/text-embedding-3-small. Sim is a similarity function as in Eq. 2 which is a cosine similarity function. Similarly, $x_{t}$ is considered member data if $s_{text}>\lambda$. This method is summarized in Algorithm 2.

### v2

@@ 2.1 Large Vision-Language Models
Large Vision-Language Models (Bai et al., 2025; Chen et al., 2024a, b; Liu et al., 2024b, a; Chen et al., 2024c; Tong et al., 2024; Zhu et al., 2025; Young et al., 2024; Zhu et al., 2023; Chen et al., 2023; Li et al., 2024c; Du et al., 2025), also known as multimodal large language models (Fu et al., 2024) are developing rapidly due to the success of language models (Wang et al., 2024; Zhao et al., 2023; Chiang et al., 2023; Touvron et al., 2023a, b; Grattafiori et al., 2024; Yang et al., 2025, 2024; Jiang et al., 2023; Team, 2023; Penedo et al., 2023; Lin et al., 2024) such as Qwen (Bai et al., 2023a; Yang et al., 2024) and LLaMA series (Touvron et al., 2023a, b). These models, like LLaVA 1.5 (Liu et al., 2024c), can generate textual outputs given textual questions and images.
@@ 2.2 Membership Inference Attack
With the development of multimodal learning, there are also some works exploring MIA on multimodal models (Hu et al., 2022b; Ko et al., 2023; Li et al., 2024b). We focus on vision-text here. Liu et al. (2021) studies MIAs on image encoder models such as CLIP vision encoder (Radford et al., 2021). It calculates the similarity scores between the augmented images’ embeddings and the scores are then used to train a classifier to infer the member status of an image. Our work is inspired by them. Ko et al. (2023) work on detecting whether an image-text pair is in the training data of CLIP models. Li et al. (2024b) investigates the single-modality MIA in LVLMs, which is a more practical scenario. They calculate the Rényi Entropy (Rényi, 1961) based on different slices of logits. Hu et al. (2025) found that member data and non-member data have different sensitivity to temperature. They then perform four attack methods under four different settings to detect whether a group of images is used in the training stage.
@@ 3.3 Intuition for Performing MIA
As we are only given the image, how to fully utilize the image to determine if the image remains a challenging question. Inspired by the findings of Liu et al. (2021) that a CLIP vision encoder tends to overfit to its training data and member images show higher similarity for two augmented versions, we assume that the image embedding of a member image from the vision encoder and modality alignment module should be more robust to image corruption than that of non-member images. The intuition is, if the model has seen the image and memorized the image, then even if the image is corrupted, e.g., some details missing, the model can still recall the details and result in embedding similar to the original one. To verify our assumption, for each image, we first apply Gaussian Blur (Bradski, 2000) as the corruption method to obtain the corrupted image, where Gaussian Blur is a smoothing technique that reduces image detail and noise by averaging pixels with a weighted Gaussian kernel. We compute the similarity score between the embedding of the original image and that of its corrupted version. The chosen model is LLaVA-1.5-7B (Liu et al., 2024a), and the dataset is the Img_Flickr dataset constructed by Li et al. (2024b). The similarity scores of member images and non-member images are shown in Figure 3. We can observe that member images normally show a higher similarity score than non-member images, which aligns well with our assumption. The scores show a clear difference between the two groups, which suggests that the similarity score between the original image embedding and its corrupted version’s embedding can be used to perform an image membership inference attack. Thus, we can utilize the robustness of member image embedding to determine its membership.
@@ 4.1 White-Box MIA via Image Similarity
where $\mathbf{e}_{i}^{v}$ is the embedding of the $i$-patch of image $\mathbf{x}_{t}$ obtained by Eq. 1 and $\mathbf{e}_{i}^{v^{\prime}}$ denotes the $i$-patch embedding of the corrupted image $\mathbf{x}_{t}^{\prime}$. $N$ is the number of patches. Sim is a similarity function. We use cosine similarity here. The larger the $s_{img}$ is, the more likely the image is viewed as a member image. The target image $\mathbf{x}_{t}$ is predicted as a member image if $s_{img}$ is bigger than a threshold $\lambda$. Algorithm 1 summarizes the image similarity-based attack.
@@ 4.2 Black-box MIA via Text Similarity
where $Emb()$ denotes a text embedding model, such as OpenAI’s text-embedding-3-small model [^1] https://platform.openai.com/docs/models/text-embedding-3-small. Sim is a similarity function as in Eq. 2 which is a cosine similarity function. Similarly, $x_{t}$ is considered member data if $s_{text}>\lambda$. This method is summarized in Algorithm 2. This attack is similar to the Target-only Inference in Hu et al. (2025) where they also use text similarity scores. The difference is that they use the average similarity score between text outputs generated by querying the model multiple times.